find the components of momentum of a 0.25 kg ball thrown with an initial velocity of 15 m/s 30
1
Expert's answer
2021-10-06T08:17:19-0400

By definition, the momentum is the velocity times mass. Thus, it is enough to find the components of velocity vector:


"v_x = 15m/s\cdot \cos30\degree =15\cdot \dfrac{\sqrt3}{2}m/s = 7.5\sqrt3m/s\\
v_y = 15m/s\cdot \sin30\degree =15\cdot \dfrac{1}{2}m/s = 7.5m/s"

The momentum is then:


"p_x = mv_x = 0.25kg\cdot7.5\sqrt3m/s = 1.1875\sqrt3\dfrac{kg\cdot m}{s}\\
p_y = mv_y = 0.25kg\cdot7.5m/s = 1.1875\dfrac{kg\cdot m}{s}\\"

Answer:

"p_x =1.1875\sqrt3\dfrac{kg\cdot m}{s}\\
p_y = 1.1875\dfrac{kg\cdot m}{s}\\"